# An Asymptotically Optimal Coordinate Descent Algorithm for Learning Bayesian Networks from Gaussian Models

### v1

@@ 5. Experiments
Setup of synthetic experiments: For all the synthetic experiments, once we specify a DAG, we generate data according to the SEM (1), where the nonzero entries of $B^{\star}$ are drawn uniformly at random from the set $\{-0.8,-0.6,0.6,0.8\}$ and diagonal entries of $\Omega^{\star}$ are chosen uniformly at random from the set $\{0.5,1,1.5\}$.
5.1. Comparison with Benchmarks
We first generate datasets from twelve publicly available networks sourced from [14] and the Bayesian Network Repository (bnlearn). These networks have different numbers of nodes, ranging from $m=6$ to $m=70$. We generate 10 independently and identically distributed datasets for each network according to the SEM described earlier with sample size $n=500$.
Table 1 compares the performance of our method CD-$\ell_{0}$ with the competing ones. First, consider small graphs ($m\leq 20$) for which the integer programming approach MICODAG achieves an optimal or near-optimal solution with a small RGAP. As expected, in terms of the accuracy of the estimated model, MICODAG tends to exhibit the best performance. For these small graphs, CD-$\ell_{0}$ performs similarly to MICODAG but attains the solutions much faster. Next, consider moderately sized graphs ($m>20$). In this case, MICODAG cannot solve these problem instances within the time limit and hence finds inaccurate models, whereas CD-$\ell_{0}$ obtains much more accurate models much faster. Finally, CD-$\ell_{0}$ outperforms GES, GSP, and CCDr-MCP in most problem instances. The improved performance of CD-$\ell_{0}$ over CCDr-MCP highlights the advantage of using $\ell_{0}$ penalization over a minimax concave penalty: $\ell_{0}$ penalization ensures that DAGs in the same Markov equivalence class have the same score, while the same property does not hold with other penalties.
Table 1. Comparison of our method, CD-$\ell_{0}$, with competing methods
$\mathrm{MICODAG}$ CCDr-MCP GES GSP CD-$\ell_{0}$
Network($m$) Time RGAP $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$
Dsep(6) $\leq 1$ 0 2.0(±$(\pm 0)$ $\leq 1$ 2.0(±$(\pm 0)$ $\leq 1$ 1.8(±$(\pm 0.6)$ $\leq 1$ 2.0(±$(\pm 0)$ $\leq 1$ 2.0(±$(\pm 0)$
Asia(8) $\leq 1$ 0 2.2(±$(\pm 0.6)$ $\leq 1$ 2.0(±$(\pm 0)$ $\leq 1$ 2.7(±$(\pm 0.9)$ $\leq 1$ 4.9(±$(\pm 1.4)$ $\leq 1$ 2.0(±$(\pm 0)$
Bowling(9) $3$ 0 2.0(±$(\pm 0)$ $\leq 1$ 4.7(±$(\pm 2.4)$ $\leq 1$ 2.4(±$(\pm 0.7)$ $\leq 1$ 5.6(±$(\pm 2.5)$ $\leq 1$ 2.2(±$(\pm 0.4)$
InsSmall(15) $\geq 750$080 7.0(±$(\pm 2.6)$ $\leq 1$ 29.9(±$(\pm 4.0)$ $\leq 1$ 24.9(±$(\pm 10.3)$ $\leq 1$ 17.2(±$(\pm 7.9)$ $\leq 1$ 8.0(±$(\pm 0)$
Rain(14) $151$ 0 2.0(±$(\pm 0)$ $\leq 1$ 9.5(±$(\pm 2.0)$ $\leq 1$ 5.4(±$(\pm 3.7)$ $\leq 1$ 17.5(±$(\pm 4.3)$ $\leq 1$ 3.3(±$(\pm 2.1)$
Cloud(16) $93$ 0 5.2(±$(\pm 0.6)$ $\leq 1$ 11.0(±$(\pm 4.1)$ $\leq 1$ 5.0(±$(\pm 1.5)$ $\leq 1$ 13.7(±$(\pm 3.0)$ $\leq 1$ 6.8(±$(\pm 2.3)$
Funnel(18) $70$ 0 2.0(±$(\pm 0)$ $\leq 1$ 2.0(±$(\pm 0)$ $\leq 1$ 4.8(±$(\pm 6.5)$ $\leq 1$ 13.0(±$(\pm 2.9)$ $\leq 1$ 2.0(±$(\pm 0)$
Galaxy(20) $237$ 0 1.0(±$(\pm 0)$ $\leq 1$ 4.6(±$(\pm 3.1)$ $\leq 1$ 1.5(±$(\pm 1.6)$ $\leq 1$ 15.8(±$(\pm 5.2)$ $\leq 1$ 1.0(±$(\pm 0)$
Insurance(27) $\geq 1350$340 22.8(±$(\pm 13.5)$ $\leq 1$ 38.4(±$(\pm 4.8)$ $\leq 1$ 30.5(±$(\pm 14.8)$ $\leq 1$ 38.5(±$(\pm 6.7)$ $\leq 1$ $(\pm 4.1)$
Factors(27) $\geq 1350$311 56.1(±$(\pm 8.4)$ $\leq 1$ 65.3(±$(\pm 7.6)$ $\leq 1$ 68.9(±$(\pm 10.5)$ $\leq 1$ 52.3(±$(\pm 7.4)$ $\leq 1$ 18.1(±$(\pm 6.7)$
Hailfinder(56) $\geq 2800$245 41.4(±$(\pm 12.6)$ $\leq 1$ 12.9(±$(\pm 3.5)$ $\leq 1$ 26.4(±$(\pm 16.2)$ $\leq 1$ 109.1(±$(\pm 10.2)$ $1.6$ 2.6(±$(\pm 1.3)$
Hepar2(70) $\geq 3500$ 5.415 76.9(±$(\pm 16.5)$ $\leq 1$ 54.6(±$(\pm 12.0)$ $\leq 1$ 71.5(±$(\pm 27.4)$ $\leq 1$ 66.3(±$(\pm 9.3)$ 11.4 5.3(±$(\pm 2.2)$
Here, MICODAG, mixed-integer convex program [25]; CCDr-MCP, minimax concave penalized estimator with coordinate descent [2]; GES, greedy equivalence search algorithm [5]; GSP, greedy sparsest permutation algorithm [19]; $d_{\mathrm{cpdag}}$, differences between the true and estimated completed partially directed acyclic graphs; RGAP, relative optimality gap. All results are computed over ten independent trials where the average $d_{\mathrm{cpdag}}$ values are presented with their standard deviations.
Large graphs: We next demonstrate the scalability of our coordinate descent algorithm for learning large DAGs with over $100$ nodes. We consider networks from the Bayesian Network Repository and generate $10$ independent datasets similar to the previous experiment. Table 2 presents the results where we see that our method CD-$\ell_{0}$ can effectively scale to large graphs and obtain better or comparable performance to competing methods, as measured by the $d_{\mathrm{cpdag}}$ metric.
Table 2. Comparison of our method, CD-$\ell_{0}$, with competing methods for large graphs
CCDr-MCP GES GSP CD-$\ell_{0}$
Network($m$) Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$
Pathfinder(109) $\leq 1$ 212.9(±$(\pm 20.7)$ $\leq 1$ 275.6(±$(\pm 16.4)$ 2.0 212.5(±$(\pm 19.5)$ 11.8 81.6(±$(\pm 16.3)$
Andes(223) $1.8$ 117.9(±$(\pm 9.6)$ $\leq 1$ 165.0(±$(\pm 28.3)$ $6.6$ 702.0(±$(\pm 42.6)$ 35.1 107.3(±$(\pm 5.9)$
Diabetes(413) $10.4$ 276.7(±$(\pm 9.7)$ $3.3$ 387.1(±$(\pm 22.2)$ $57.8$ 1399.8(±$(\pm 19.1)$ 881.9 286.6(±$(\pm 15.9)$
See Table01 for the description of the methods. All results are computed over ten independent trials where the average $d_{\mathrm{cpdag}}$ values are presented with their standard deviations.
5.2. Convergence of CD-$\ell_{0}$ Solution to An Optimal Solution
@@
Theorem 7 states that as the sample size tends to infinity, CD-$\ell_{0}$ identifies an optimally scoring model. To see how fast the asymptotic kicks in, we generate three synthetic DAGs with $m=10$ nodes where the total number of edges is chosen from the set $\{7,12,21\}$. We obtain $10$ independently and identically distributed datasets according to the SEM described earlier with sample size $n=\{50,100,200,300,400,500\}$. In Figure 2(left, middle), we compute the normalized difference $(\mathrm{obj}^{\mathrm{method}}-\mathrm{obj}^{\mathrm{opt}})/\mathrm{obj}^{\mathrm{opt}}$ as a function of $n$ for the three graphs, averaged across the ten independent trials. Here, $\mathrm{obj}^{\mathrm{method}}$ is the objective value obtained by the corresponding method (CD-$\ell_{0}$ or GES), while $\mathrm{obj}^{\mathrm{opt}}$ is the optimal objective obtained by the integer programming approach MICODAG. For moderately large sample sizes (e.g., $n=200$), CD-$\ell_{0}$ attains the optimal objective value, whereas GES does not. In Figure 2 (right), for the graph with 21 arcs, we see that CD-$\ell_{0}$ can achieve the same accuracy while being computationally much faster to solve.
Figure 2. Convergence of CD-$\ell_{0}$ to an optimal solution
@@
Left: normalized difference, as a function of sample size $n$, between the optimal objective value of (3) found using the integer programming approach MICODAG and the objective value obtained by CD-$\ell_{0}$ for three different graphs; Middle: normalized difference of objectives of solutions obtained from MICODAG and GES; Right: comparison of computational cost of CD-$\ell_{0}$, MICODAG, and GES for the DAG with $21$ edges. All results are computed and averaged over ten independent trials.
5.3. Real Data from Causal Chambers

### v3

@@ 6 Synthetic and Real Experiments
Setup of synthetic experiments: For all the synthetic experiments, following the experimental setup in Xu et al. (2024), once we specify a DAG, we generate data according to the SEM (1), where the nonzero entries of $B^{\star}$ are drawn uniformly at random from the set $\{-0.8,-0.6,0.6,0.8\}$. Unless otherwise specified, the diagonal entries of $\Omega^{\star}$ are drawn uniformly at random from the set $\{0.8,1,1.2\}$. This generation of noise variance introduces some heteroscedasticity. Later, we vary this set to study how our method performs under different levels of heteroscedasticity.
6.1 Convergence of CD-$\ell_{0}$ solution to an optimal solution
@@
Theorem 7 states that as the sample size tends to infinity, CD-$\ell_{0}$ identifies an optimally scoring model. To observe how quickly the asymptotics take effect, we generate three synthetic DAGs with $m=10$ nodes, where the total number of edges is chosen from the set $\{7,12,21\}$. We obtain $10$ independently and identically distributed data sets according to the SEM described earlier with sample size $n=\{100,200,300,400,500,800,1000,1600,3200\}$. In Figure 2, we compute the normalized difference $(\mathrm{obj}^{\mathrm{method}}-\mathrm{obj}^{\mathrm{opt}})/\mathrm{obj}^{\mathrm{opt}}$ as a function of $n$ for the three graphs, averaged across the ten independent trials. Here, $\mathrm{obj}^{\mathrm{method}}$ is the objective value obtained by the corresponding method (CD-$\ell_{0}$ or GES), while $\mathrm{obj}^{\mathrm{opt}}$ is the optimal objective of (3) obtained by the integer programming approach MICODAG. For moderately large sample sizes, CD-$\ell_{0}$ attains the optimal objective value, whereas GES does not. Meanwhile, CD-$\ell_{0}$ obtains a near-optimal solution at about 0.1% of time, as compared to MICODAG. In the left panel of Figure 2, we run CD-$\ell_{0}$ with three different random orderings. In all cases, the objective converges to zero as expected, though at varying rates. The worse-case rate is $\mathcal{O}(\sqrt{d_{\mathrm{max}}^{2}m^{4}\log m/n})$ according to Theorem 7.
Figure 2: Convergence of CD-$\ell_{0}$ to an optimal solution
@@
Left: normalized difference of CD-$\ell_{0}$ with 3 different random orderings; Middle: normalized difference for different graphs; Right: normalized difference of objectives of solutions obtained from MICODAG and GES. All results are computed and averaged over ten independent trials.
6.2 Comparison to benchmarks under near-homoscedastic error
We next generate data sets from twelve publicly available networks sourced from Manzour et al. (2021) and the Bayesian Network Repository (bnlearn). These networks have different numbers of nodes, ranging from $m=6$ to $m=70$. We generate 10 independently and identically distributed data sets for each network according to the SEM described earlier with sample size $n=500$. Since the diagonal entries of $\Omega^{\star}$ are close (with the largest deviation being $0.8$ to $1.2$), our algorithm performs very similarly to TD, since the topological ordering obtained by Algorithm 3 is likely to be correct in this near-homoscedastic case, and both methods perform well. For this reason, we present the results for the nearly homoscedastic case in Appendix C, and instead focus on a slightly higher degree of heteroscedasticity to highlight that our method is more robust to heteroscedasticity than TD. To that end, the diagonal entries of $\Omega^{\star}$ are drawn uniformly from $\{0.6,1,1.2\}$.
Table 1 compares the performance of our method CD-$\ell_{0}$ with the competing ones. First, consider small graphs ($m\leq 20$) for which the integer programming approach MICODAG achieves an optimal or near-optimal solution with a small RGAP. As expected, in terms of the accuracy of the estimated model, MICODAG exhibits the best performance. For these small graphs, CD-$\ell_{0}$ performs similarly to MICODAG but attains the solutions much faster. Next, consider moderately sized graphs ($m>20$). In this case, MICODAG cannot solve these problem instances within the time limit and hence finds inaccurate models, whereas CD-$\ell_{0}$ obtains much more accurate models much faster. Finally, CD-$\ell_{0}$ outperforms CCDr-MCP, GES, TD, and NOTEARS in most problem instances. The improved performance of CD-$\ell_{0}$ over CCDr-MCP highlights the advantage of using $\ell_{0}$ penalization over a minimax concave penalty: $\ell_{0}$ penalization ensures that DAGs in the same Markov equivalence class have the same score. In contrast, this property does not hold for other penalties. Our method outperforms NOTEARS in terms of the $d_{\mathrm{cpdag}}$ for most graphs. Both methods are designed for computational efficiency, but NOTEARS uses a least squares loss, which implicitly assumes homoscedastic (equal variance) noise (van de Geer and Bühlmann, 2013). In contrast, our approach directly optimizes the original likelihood function, allowing us to account for heteroscedasticity. This results in improved accuracy and robustness under varying noise conditions.
Table 1: Comparison of our method, CD-$\ell_{0}$, with competing methods
$\mathrm{MICODAG}$ CCDr-MCP GES TD NOTEARS CD-$\ell_{0}$
Network($m$) Time RGAP $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$
Dsep(6) $\leq 1$ 0 2.0(±0)($\pm$ 0) $\leq 1$ 2.0(±0)($\pm$ 0) $\leq 1$ 2.0(±0)($\pm$ 0) $\leq 1$ 2.1(±0.3)($\pm$ 0.3) $\leq 1$ 2.0(±0.0)($\pm$ 0.0) $\leq 1$ 2.0(±0)($\pm$ 0)
Asia(8) $\leq 1$ 0 2.0(±0)($\pm$ 0) $\leq 1$ 2.3(±0.9)($\pm$ 0.9) $\leq 1$ 2.1(±0.3)($\pm$ 0.3) $\leq 1$ 2.1(±0.3)($\pm$ 0.3) $\leq 1$ 3.5(±2.4)($\pm$ 2.4) $\leq 1$ 2.1(±0.3)($\pm$ 0.3)
Bowling(9) 1.2 0 2.0(±0)($\pm$ 0) $\leq 1$ 3.1(±1.9)($\pm$ 1.9) $\leq 1$ 11.8(±8.0)($\pm$ 8.0) $\leq 1$ 2.2(±0.6)($\pm$ 0.6) $\leq 1$ 3.6(±0.8)($\pm$ 0.8) $\leq 1$ 2.0(±0)($\pm$ 0)
InsSmall(15) $\geq 750$ .007 7.2(±1.0)($\pm$ 1.0) $\leq 1$ 27.0(±4.1)($\pm$ 4.1) $\leq 1$ 37.2(±4.7)($\pm$ 4.7) $\leq 1$ 3.9(±4.8)($\pm$ 4.8) $\leq 1$ 10.9(±1.6)($\pm$ 1.6) $\leq 1$ 4.0(±4.6)($\pm$ 4.6)
Rain(14) 56.0 0 2.0(±0)($\pm$ 0) $\leq 1$ 8.5(±2.1)($\pm$ 2.1) $\leq 1$ 15.5(±2.2)($\pm$ 2.2) $\leq 1$ 6.6(±2.1)($\pm$ 2.1) $\leq 1$ 5.4(±2.1)($\pm$ 2.1) $\leq 1$ 5.5(±2.0)($\pm$ 2.0)
Cloud(16) 9.7 0 12.7(±5.5)($\pm$ 5.5) $\leq 1$ 8.5(±4.5)($\pm$ 4.5) $\leq 1$ 10.6(±8.2)($\pm$ 8.2) $\leq 1$ 11.9(±7.8)($\pm$ 7.8) $\leq 1$ 17.6(±3.7)($\pm$ 3.7) $\leq 1$ 11.6(±6.5)($\pm$ 6.5)
Funnel(18) 8.0 0 2.0(±0)($\pm$ 0) $\leq 1$ 3.3(±1.7)($\pm$ 1.7) $\leq 1$ 2.6(±1.3)($\pm$ 1.3) $\leq 1$ 6.4(±2.5)($\pm$ 2.5) $\leq 1$ 2.9(±1.4)($\pm$ 1.4) $\leq 1$ 5.4(±2.0)($\pm$ 2.0)
Galaxy(20) 48.6 0 1.0(±0)($\pm$ 0) $\leq 1$ 9.1(±4.3)($\pm$ 4.3) $\leq 1$ 26.2(±0.6)($\pm$ 0.6) $\leq 1$ 11.7(±6.7)($\pm$ 6.7) $\leq 1$ 12.4(±2.1)($\pm$ 2.1) $\leq 1$ 9.7(±4.9)($\pm$ 4.9)
Insurance(27) $\geq 1350$ .272 18.5(±8.5)($\pm$ 8.5) $\leq 1$ 32.5(±6.0)($\pm$ 6.0) $\leq 1$ 24.2(±7.9)($\pm$ 7.9) $\leq 1$ 16.8(±2.2)($\pm$ 2.2) $\leq 1$ 12.8(±4.0)($\pm$ 4.0) $\leq 1$ 18.3(±1.9)($\pm$ 1.9)
Factors(27) $\geq 1350$ .242 60.4(±7.3)($\pm$ 7.3) $\leq 1$ 65.1(±6.4)($\pm$ 6.4) $\leq 1$ 68.5(±9.3)($\pm$ 9.3) $\leq 1$ 26.9(±6.4)($\pm$ 6.4) $\leq 1$ 41.7(±5.1)($\pm$ 5.1) $\leq 1$ 25.2(±9.7)($\pm$ 9.7)
Hfinder(56) $\geq 2800$ .181 12.7(±4.6)($\pm$ 4.6) $\leq 1$ 14.9(±3.8)($\pm$ 3.8) $\leq 1$ 27.5(±14.4)($\pm$ 14.4) 1.2 38.3(±7.7)($\pm$ 7.7) 1.1 17.3(±4.0)($\pm$ 4.0) $\leq 1$ 45.1(±8.9)($\pm$ 8.9)
Hepar(70) $\geq 3500$ .868 45.6(±9.8)($\pm$ 9.8) $\leq 1$ 51.6(±9.5)($\pm$ 9.5) $\leq 1$ 71.9(±19.7)($\pm$ 19.7) 2.6 54.2(±9.8)($\pm$ 9.8) 2.2 43.9(±5.4)($\pm$ 5.4) $\leq 1$ 38.5(±5.9)($\pm$ 5.9)
Here, MICODAG, mixed-integer convex program (Xu et al., 2024); CCDr-MCP, minimax concave penalized estimator with coordinate descent (Aragam et al., 2019); GES, greedy equivalence search algorithm (Chickering, 2002); TD, top-down method (Chen et al., 2019); NOTEARS, (Zheng et al., 2018); $d_{\mathrm{cpdag}}$, differences between the true and estimated completed partially directed acyclic graphs; RGAP, relative optimality gap. All results are computed over ten independent trials, where the average $d_{\mathrm{cpdag}}$ values are presented with their standard deviations.
Large graphs: We next demonstrate the scalability of our coordinate descent algorithm for learning large DAGs with over $100$ nodes. We consider networks from the Bayesian Network Repository and generate $10$ independent datasets similar to those in the previous experiment. The method TD is excluded since it cannot scale to very large graphs. Table 2 presents the results where we see that our method CD-$\ell_{0}$ can effectively scale to large graphs and obtain better or comparable performance to competing methods, as measured by the $d_{\mathrm{cpdag}}$ metric.
Table 2: Comparison of our method, CD-$\ell_{0}$, with competing methods for large graphs
CCDr-MCP GES CD-$\ell_{0}$
Network($m$) Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$ Time $d_{\mathrm{cpdag}}$
Pathfinder(109) $5.5$ 206.6(±4.86)($\pm$ 4.86) $\leq 1$ 253.8(±19.52)($\pm$ 19.52) 24.0 95.0(±18.40)($\pm$ 18.40)
Andes(223) $12.6$ 121.5(±5.23)($\pm$ 5.23) $1.9$ 152.8(±22.26)($\pm$ 22.26) 108.9 98.4(±6.82)($\pm$ 6.82)
Diabetes(413) $42.6$ 322.9(±9.39)($\pm$ 9.39) $10.7$ 458.7(±28.73)($\pm$ 28.73) 732.3 158.4(±10.76)($\pm$ 10.76)
See Table01 for the description of the methods. All results are computed over ten independent trials, where the average $d_{\mathrm{cpdag}}$ values are presented with their standard deviations.
6.3 Severe heteroscedastic error and limitations of our method
One limitation of our algorithm is that under severe heteroscedasticity, recovering an accurate topological ordering becomes difficult, which can degrade performance. To demonstrate this, we draw diagonal entries of $\Omega^{\star}$ uniformly at random from the set $\{0.2,1,5\}$, so that the noise distribution is highly heteroscedastic. In this setup, Assumption 3 is unlikely to hold, making the input topological ordering less reliable for both TD and our method. Despite this, the results in Table 3 show that both approaches remain competitive.
Table 3: Comparison of GES, Top-down (TD), and CD-$\ell_{0}$ using datasets with very large heteroscedasticity.
$d_{\mathrm{cpdag}}$
Network($m$) GES TD $\text{CD-}{\ell_{0}}$
Dsep(6) $4.9(\pm 1.4)$ $5.4(\pm 0.8)$ $4.7(\pm 1.2)$
Asia(8) $6.4(\pm 2.8)$ $14.4(\pm 1.6)$ $16.4(\pm 1.0)$
Bowling(9) $8.2(\pm 6.8)$ $10.0(\pm 0)$ $10.0(\pm 0)$
InsuranceSmall(15) $22.5(\pm 5.1)$ $18.7(\pm 2.5)$ $17.2(\pm 1.6)$
Rain(14) $26.6(\pm 1.8)$ $10.5(\pm 2.8)$ $9.9(\pm 2.3)$
Cloud(16) $14.9(\pm 3.6)$ $19.4(\pm 1.8)$ $19.0(\pm 3.8)$
Funnel(18) $10.0(\pm 2.1)$ $9.4(\pm 2.1)$ $7.6(\pm 1.4)$
Galaxy(20) $6.0(\pm 2.4)$ $35.6(\pm 3.9)$ $26.5(\pm 1.6)$
Insurance(27) $61.1(\pm 7.7)$ $65.8(\pm 3.6)$ $46.9(\pm 4.1)$
Factors(27) $74.1(\pm 4.8)$ $47.5(\pm 7.5)$ $40.0(\pm 10.0)$
Hfinder(56) $41.0(\pm 5.1)$ $93.6(\pm 2.8)$ $100.0(\pm 3.7)$
Hepar(70) $68.5(\pm 11.9)$ $124.1(\pm 8.7)$ $91.5(\pm 12.5)$
Our algorithm CD-$\ell_{0}$ does not rely on a specific topological ordering and can take any ordering as input—not just the one produced by TD. In practice, if an accurate method for obtaining the topological ordering in a heteroscedastic setting is available, it can be used as input to Algorithm 1.
6.4 Real data from causal chambers